PLEASE ANSWER E: Ceramics, Inc. (CI), a company that manufactures bath tiles, is interested in measuring inventory effectiveness. Last year the cost of goods sold at CI was $200,000 on sales of $250,000. The average inventory in dollars was $12,000. At the end of the year, CI had outstanding receivables of $30,000 and payables of $20,000.

a) Calculate the inventory turnover for CI.

b) Calculate the days of inventory. Assume that CI operates 6 days per week and 52 weeks per year.

c) Calculate the weeks of inventory. Assume 52 weeks per year. d) What is the cash-to-cash cycle time at CI?

e) Assume it is December 31, and CI has exactly $12,000 in inventory. CI has a forecast of $11,000 in sales for January and $9,000 in February. How many days of forecast sales can be met with the current inventory given there are 23 work days in January and 21 in February? (Use specifics, not averages.)
Answer #1

a) Calculate the inventory turnover for CI.

COGS = $ 200,000

Average Inventory = $ 12,000

Inventory turnover = COGS / Average inventory

= 200000 / 12000

= 16.67

b) Calculate the days of inventory. Assume that CI operates 6 days per week and 52 weeks per year.

Number of days per year = 52 weeks * 6 weeks = 312 days

Days of inventory , DIO = 312 * Average inventory / COGS

= 312 * 12000 / 200000

= 18.72 days

c) Calculate the weeks of inventory. Assume 52 weeks per year.

Weeks of Inventory = 18.72/6

= 3.12 weeks

d) What is the cash-to-cash cycle time at CI?

Days of Sales outstanding (DSO) = d * AR / Sales

= 312*30000/250000

= 37.44 days

Days of Payable outstanding, DPO = d*AP/COGS

= 312 * 20000 / 200000

= 31.2 days

Cash to cash cycle time = DIO + DSO - DPO

= 18.72 + 37.44 - 31.20

= 24.96 days

e) Assume it is December 31, and CI has exactly $12,000 in inventory. CI has a forecast of $11,000 in sales for January and $9,000 in February. How many days of forecast sales can be met with the current inventory given there are 23 work days in January and 21 in February?

The entire January forcarsted sales can be met, as ithe forcasted sales are less than the inventory on 31st december.

At the end of january,

Inventory left = 12000 - 11000 =1000

Forcated salea of Feb = 9000 for 21 days.

Therefore 1 day = 9000 / 21= 428.57

Let the number of days in feb where the demand will be met be x.

Therefore, 428.57x = 1000

x = 1000/428.57= 2.3 days, approximately 2 days

Total number of days forcasted sales can be met by cirrent inventory = 23 + 2.3 = 25.3 days,

approximately 25 days